# 1in1in1in1in

### v9

@@ 5.2 Open contest at MLCAD 2023
The EDA industry and research community regularly organize open research contests to keep track of SOTA on industry-produced chip designs, to evaluate published methods and codes, and to perform fair comparisons. A recent contest is relevant to our meta-analysis.
In 2023, the IEEE/ACM Workshop on Machine Learning in Computer-Aided Design (MLCAD) held an open research contest for macro placement for FPGAs and ASICs ( https://mlcad-workshop.org/1st-mlcad-contest/). The contest followed the playbook established at other CAD conferences such as DAC, ICCAD, and ISPD, and particularly successful in evaluating algorithmic techniques for physical design since the ISPD 2005 placement contest [30, 37, 38, 39, 40, 42]:
  • a challenge problem is announced along with a research infrastructure that includes chip design examples and metrics;
  • several months after sign-up, final submissions are collected and evaluated by the organizers on hidden design examples (benchmarks); the same computational resources are used for all contestants to ensure fairness;
@@
Despite technical differences from infrastructure in [1], the contest offered a great opportunity for Google and the authors of [1] to showcase the versatility and quality of their RL technique. [^12] RL was advertised in [1] as a general technique for combinatorial optimization, so should be able to handle various macro placement formulations. The contest attracted 19 participants, of them 8 finalists — university teams from Taiwan, Hong Kong, China, Canada and the US (several students per team). Google did not participate.131313 Google did not skip a similar research contest at the Intl. Workshop on Logic Synthesis (IWLS 2023). That contest ( https://github.com/alanminko/iwls2023-ls-contest) focused on chip logic design rather than physical design (two nearby fields). Google won the first place using long-running parallel Simulated Annealing. The winning team overlapped with Google Team 1 working on macro placement: Sergio Guadarrama (quoted in [67]), the senior staff software engineer co-authored the ISPD 2022 paper [5] with the lead authors of [1] and then tweeted on 5/3/22 that he and his team at Google “independently replicated” the work in [1].
Contest results were announced on September 13, 2023. According to the participants’ own descriptions, top six teams used traditional analytical optimization methods sans ML — based on either DREAMPlace [50] (derived from RePlAce) or SimPL [35]. The absence of RL solutions was noteworthy, given that Google “open-sourced” the methods of [1] in [4]. Evaluation scores on hidden benchmarks generally mirrored the trends seen on public benchmarks, which is not surprising in the absence of ML methods (that sometimes overfit to training data). Overall contest results were consistent with the conclusions in [6, 8] that traditional optimization methods in circuit placement produced better macro placement results than RL from [1] did, and finished faster. The results in [8] complete the overall picture by evaluating chip metrics used in [1]: power, circuit delay, and the total area of standard cells. Industry chips are traditionally evaluated using the PPA triad (Power, delay Performance, and chip Area), but macro placement methodologies considered do not alter chip area.
@@ 6 Rebuttals to critiques of [1]
Despite critical media coverage [54, 55, 61, 65] and technical questions raised, the authors of [1, 4] failed to remove remaining obstacles to reproducibility [64] of the methods and results in [1]. The UCSD team’s engineering effort overcame those obstacles, and they followed up on the work of Google Team 2 [6] that criticized [1], then analyzed many issues listed in Sections 3 and 4. Google Team 2 had access to Google TPU designs and the source code used in [1] before the CT GitHub repo [4] appeared. The UCSD authors of [7, 8] had access to Circuit Training (CT) [4] and benefited from a lengthy involvement of Google Team 1 engineers, but not access to SA code used in [1, 6] or other key pieces of code missing from [4]. Yet, the results in [6] and [7, 8] corroborate each other, and their qualitative conclusions are consistent. UCSD results for Ariane-NG45 closely match those by Google Team 1 engineers, and [8, Figure 4] shows that CT training curves of Ariane-NG45 generated at UCSD match those produced by Google Team 1 engineers. Google Team 1 engineers carefully reviewed the paper [8] and the work in Fall 2022 and Winter 2023, raising no objections [7, FAQ].
@@
Old benchmarks. Another objection [65] is that public benchmarks [23] used in [6, 8] allegedly use outdated infrastructure. But those circuits have been evaluated with the HPWL objective, which scales accurately under geometric 2D scaling of chip designs and remains appropriate for all technology nodes (Section 2). Per [2], ICCAD benchmarks were requested [2] by Peer Reviewer #3 of [1]. When [6, 8] implemented this ask, Google RL ran into trouble before routing became relevant: it lost by 20% or so in HPWL optimization (HPWL is the simplest yet most important term of the proxy cost optimized by CT/RL [1, 4]).
Not training until convergence in experiments in [8]. This concern was promptly addressed in FAQ #15 in [7]: “ ‘training until convergence’ is not described in any of the guidelines provided by the CT GitHub repo for reproducing the results in the Nature paper.” The authors of [8] followed guidelines by Google in [4]. Later, their additional experiments indicated that “training until convergence worsens some key chip metrics while improving others, highlighting the poor correlation between proxy cost and chip metrics. Overall, training until convergence does not qualitatively change comparisons to results of Simulated Annealing and human macro placements reported in the ISPD 2023 paper.” RL-vs-SA experiments in [6] predated [4], so trained until convergence with the 6-hour protocol from [1].
@@ 7 Can the work in [1] be used?
The Nature paper [1] claimed applications to recent Google TPU chips, providing credence to the notion that those methods improved State of the Art. But aside from vague general claims, no chip-metric improvements were reported for specific production chips.141414[1, Table 1] shows results for TPU designs of an earlier generation (that is, on chips that were already manufactured at the time). Assuming substantial use in production, more recent TPU design blocks must have used [1, 4] for tape-out. Section 5 shows that the methods of [1, 4] lag behind SOTA, e.g., Simulated Annealing from the 1980s [11, 13, 14, 15]. Moreover, a strong Google-internal implementation of SA from [6] could serve as a drop-in replacement of RL in [1, 4]. Without inside knowledge of details, we speculate how to reconcile the claimed use in TPUs with Google CT/RL lagging behind SOTA (per [6, 8]).
  • Given the high variance of RL results in terms of chip-timing metrics TNS and WNS (due to low correlation with the proxy metric), trying many independent randomized attempts with variant proxy cost functions and hyperparameter settings may improve best-seen results [59], increasing runtime by many times over. But SA can also be used this way.
  • Using in-house methods, even if inferior, is a common methodology in industry practice called dogfooding (“eat your own dogfood”). In most chip designs, some blocks are not critical (do not affect chip speed) and are good candidates for dogfooding. This can explain selective “production use” and reporting.
  • Per [1], the results of RL were postprocessed by SA but the CT FAQ [4] disclaimed this postprocessing — postprocessing was used in the TPU design flow but not when comparing RL to SA. But since full-fledged SA consistently beats RL [6, 8], SA could substitute for RL (initial locations can be accommodated using an adaptive temperature schedule in SA).
  • Google Team 1’s follow-up [5] shows (in Figure 7) that pre-training improves results only when pre-training on essentially the same design. Perhaps, Google is leveraging RL when performing multiple revisions to IC designs — a valid context, but not described in [1]. Moreover, commercial EDA tools are orders of magnitude faster when running from scratch [8], so pre-training does not close the gap with [1].
@@
    RL’s placement of macros on a coarse grid limits their locations (Figure 1). When a human designer was told to ignore the course grid, they found better macro locations [8]. Commercial EDA tools also avoid this limitation and outperform Google CT/RL.
@@ 8 Conclusions
This meta-analysis discusses the reproduction and evaluation of results in the Nature paper [1], as well as the validity of methods, results and claims in that paper. Based on crosschecked newer data, we draw conclusions with ample redundancy (resistant to isolated mistakes): the integrity of [1] is substantially undermined owing to errors in the conduct, analysis and reporting of its study. Omissions, inconsistencies, mistakes, and misrepresentations impacted methods, data, results and interpretation in [1]. Table 4 makes it clear that serious issues were raised at Google about the work many months before the Nature publication [1].
@@ 8.1 Conclusions about [1]
We crosscheck the results reported in [6, 7, 8] and also account for [2, 4, 5, 51], then summarize conclusions drawn from these works. This confirms many of the initial doubts about [1] and identifies additional deficiencies. As a result, it is clear that [1] is misleading in several ways, such that the readers can have no confidence in the top-line claims of [1]. nor its conclusions. [1] did not improve SOTA while the methods and results of the original paper were not reproducible from the descriptions provided, contrary to stated editorial policies at Nature (see Section 8.3). The reliance on proprietary TPU designs for evaluation, along with insufficient reporting of experiments, continues to obstruct reproducibility of the methods and the results. The authors of [6] had access to Google internal code whereas [8] reverse-engineered and/or reimplemented missing components. Google Team 2 and the UCSD team drew consistent conclusions from similar experiments, and each team made additional observations.
@@ 8.2 Conclusions for chip-design tech
    Machine-learning from entire chip designs is hard: learning from diverse designs might only improve runtime and not quality of results on any one design [5]. Learning from earlier versions of the same design can be useful in some cases, but should be compared to warm-starting Simulated Annealing with an initial placement and adaptive temperature schedule.
@@ 8.3 Policy implications
Theoretical arguments and empirical evidence suggest that numerous published papers across various fields do not replicate and are likely false [10, 9, 41, 75, 82]. The developments discussed in this meta-analysis add to the so-called reproducibility crisis that undermines trust in published research results [70, 75]. In response to this crisis, some demand “reproduce or it didn’t happen” [79].
Retraction Watch now tracks 5000 retractions per year, including prominent cases of research misconduct [73, 75]. Per [70], “research misconduct is a serious problem and (probably) getting worse”, which makes it even more important to separate honest mistakes from deliberate exaggerations and misconduct [71, 72, 73, 76, 82]. To this end, see Table 4. Institutional response is needed, and opportunities for future reforms are discussed in [69, 76, 83]. Here we make a modest contribution to this far-reaching discussion by making specific suggestions.
@@
It is unclear why Google did not allow publishing [6] (coauthored by the author of this meta-analysis), even after its results and conclusions were corroborated by the published paper [8] written at UCSD with lengthy involvement from Google. Granted, [6] and [7, 8] found major flaws in [1], but “a commitment to open inquiry, intellectual rigor, integrity, and collaboration” must protect legitimate research, even if it is politically inconvenient.
@@
Specifically for [1], the Nature editorial [3] insisted that “the technical expertise must be shared widely.” Furthermore, when manuscript authors neglect requests for public benchmarking and obstruct reproducibility, their technical claims should be viewed with suspicion [64] (especially if they later disagree with comparisons to their work [60]). Per peer review file [2], the acceptance of the Nature paper was conditional on the release of code and data in the second revision, but this did not happen when the paper was published or later, per [8]. The Nature paper [1] was amended by the authors to claim that the code had been made available (see the “Data and Code Availability” disclaimer). But serious omissions remain in the released code. This is particularly concerning because ($i$) [1] omitted key comparisons and details, and ($ii$) fraud was alleged under oath in a California court by a Google whistleblower who worked on the project [59]. This makes reproducibility more critical.
@@ 8.4 Nature editors investigate
Nine months later, the investigation has not produced results. And we do not have high hopes, given the Nature News team’s detailed account of an unrelated scandal involving two retracted Nature papers on high-temperature superconductivity by Ranga Dias [80, 81].
The retraction does not state what Hamlin and Ramshaw found in the post-publication review process instigated by Nature: that the raw data were probably fabricated.
Per Nature editor Karl Ziemelis, “Allegations of possible misconduct are outside the remit of peer review and more appropriately investigated by the host institution.” That works poorly: three Univ. of Rochester investigations “did not find evidence of misconduct,” and only the fourth found “data reliability concerns.”
Publicly, Dias continued to insist that CSH was legitimate and that the retraction was simply down to an obscure technical disagreement.
Company-internal investigations have even fewer incentives to disclose misconduct to the public, whereas litigation is usually resolved with private settlements and non-disclosure agreements to avoid public admission of wrongdoing.

### v10

@@ 5.2 Open contest at MLCAD 2023
The EDA industry and research community regularly organize open research contests to keep track of SOTA on industry-produced chip designs, to evaluate published methods and codes, and to perform fair comparisons. A recent contest is relevant to our meta-analysis. In 2023, the IEEE/ACM Workshop on Machine Learning in Computer-Aided Design (MLCAD) held an open research contest for macro placement for FPGAs and ASICs ( https://mlcad-workshop.org/1st-mlcad-contest/). The contest followed the playbook established at other CAD conferences such as DAC, ICCAD, and ISPD, and particularly successful in evaluating algorithmic techniques for physical design since the ISPD 2005 placement contest [30, 37, 38, 39, 40, 42]:
  • a challenge problem is announced along with chip design examples and metrics;
  • several months after sign-up, final submissions are collected and evaluated by the organizers on hidden design examples (benchmarks); the same compute resources are available to all contestants (fairness);
@@
Despite technical differences from infrastructure in [1], the contest offered a great opportunity for Google and the authors of [1] to showcase the versatility and quality of their RL technique. [^12] RL was pitched in [1] as a general technique for combinatorial optimization, so must handle various macro-placement tasks. The contest attracted 19 participants, of them 8 finalists — academic teams from Taiwan, Hong Kong, China, Canada and the US (several students per team). Google did not take part.131313 Google did not skip a similar research contest at the Intl. Workshop on Logic Synthesis (IWLS 2023). That contest ( https://github.com/alanminko/iwls2023-ls-contest) focused on chip logic design rather than physical design (two nearby fields). Google won the first place using long-running parallel Simulated Annealing (but during the IWLS 2024 Contest another team outperformed Google results without using ML or DNNs, per https://github.com/alanminko/iwls2023-ls-contest). The winning team overlapped with Google Team 1 working on macro placement: Sergio Guadarrama (quoted in [69]), the senior staff software engineer co-authored the ISPD 2022 paper [5] with the lead authors of [1] and then tweeted on 5/3/22 that he and his team at Google “independently replicated” the work in [1].
Contest results were announced on September 13, 2023. According to the participants’ own descriptions, top six teams used traditional analytical optimization methods sans ML — based on either DREAMPlace [51] (derived from RePlAce) or SimPL [35]. The absence of RL solutions was noteworthy, given that Google “open-sourced” the methods of [1] in [4]. Evaluation scores on hidden benchmarks generally mirrored the trends seen on public benchmarks, which is not surprising in the absence of ML methods (that sometimes overfit to training data). Overall contest results were consistent with the conclusions in [6, 8] that traditional optimization methods in circuit placement produced better macro placement results than RL from [1] did, and finished faster. The results in [8] complete the overall picture by evaluating chip metrics used in [1]: power, circuit delay, and the total area of standard cells. Industry chips are typically evaluated using the PPA triad (Power, delay Performance, and chip Area), but macro placement methodologies considered do not alter chip area.
@@ 6 Rebuttals to critiques of [1]
Despite critical media coverage [55, 56, 63, 67] and technical questions raised, the authors of [1, 4] failed to remove remaining obstacles to reproducibility [66] of the methods and results in [1]. The UCSD team’s engineering effort overcame those obstacles, and they followed up on the work of Google Team 2 [6] that criticized [1], then analyzed many issues listed in Sections 3 and 4. Google Team 2 had access to Google TPU designs and the source code used in [1] before the CT GitHub repo [4] appeared. The UCSD authors of [7, 8] had access to Circuit Training (CT) [4] and benefited from a lengthy involvement of Google Team 1 engineers, but had no access to SA code used in [1, 6] or other key pieces of code missing from [4]. Yet, the results in [6] and [7, 8] corroborate each other, and their qualitative conclusions are consistent. UCSD results for Ariane-NG45 closely match those by Google Team 1 engineers, and [8, Figure 4] shows that CT training curves of Ariane-NG45 generated at UCSD match those produced by Google Team 1 engineers. Google Team 1 engineers carefully reviewed the paper [8] and the work in Fall 2022 and Winter 2023, raising no objections [7, FAQ].
@@
Old benchmarks. Another objection [67] is that public benchmarks [23] used in [6, 8] allegedly use outdated infrastructure. But those circuits have been evaluated with the HPWL objective, which scales accurately under geometric 2D scaling of chip designs and remains appropriate for all technology nodes (Section 2). Per [2], ICCAD benchmarks were requested [2] by Peer Reviewer #3 of [1]. When [6, 8] implemented this ask, Google RL ran into trouble before routing became relevant: RL lost by 20% or so in HPWL optimization (HPWL is the simplest yet most important term of the proxy cost optimized by CT/RL [1, 4]).
Not training until convergence in experiments in [8]. This concern was promptly addressed in FAQ #15 in [7]: “ ‘training until convergence’ is not described in any of the guidelines provided by the CT GitHub repo for reproducing the results in the Nature paper.” The authors of [8] followed Google’s guidelines in the CT repo [4]. Later, their additional experiments indicated that “training until convergence worsens some key chip metrics while improving others, highlighting the poor correlation between proxy cost and chip metrics. Overall, training until convergence does not qualitatively change comparisons to results of Simulated Annealing and human macro placements reported in the ISPD 2023 paper.” RL-vs.-SA experiments in [6] predated [4], so trained until convergence per the six-hour protocol from [1].
@@ 7 Can the work in [1] be used?
The Nature paper [1] claimed applications to recent Google TPU chips, providing credence to the notion that those methods improved State of the Art. But aside from vague general claims, no chip-metric improvements were reported for specific production chips.141414[1, Table 1] shows results for TPU designs of an earlier generation (that is, on chips that were already manufactured at the time). Assuming substantial use in production, more recent TPU design blocks must have used [1, 4] for tape-out. Section 5 shows that the methods of [1, 4] lag behind Simulated Annealing from the 1980s [11, 13, 14, 15]. Moreover, the Google-internal implementation of SA from [6] (or the public one from [8, 7]) could serve as a drop-in replacement of RL in [1, 4]. Without inside knowledge, we speculate how to reconcile the claimed use in TPUs with Google CT/RL lagging behind SOTA (per [6, 8]).
  • Given the high variance of chip-timing metrics TNS and WNS in RL results (due to low correlation with the proxy metric), trying many independent randomized attempts with variant proxy cost functions and hyperparameter settings may improve best-seen results [61], with much greater runtimes. But SA can also be used this way.
  • Using in-house methods, even if inferior, is a common methodology in industry practice called dogfooding (“eat your own dogfood”). Blocks that are not critical (do not affect chip speed) are good candidates for dogfooding. This can explain selective “production use” and reporting.
  • Per [1], the results of RL were postprocessed by SA but the CT FAQ [4] disclaimed this postprocessing — postprocessing was used in the TPU design flow but not when comparing RL to SA. But since full-fledged SA consistently beats RL [6, 8], SA could substitute for RL (initial locations can be used with adaptive temperature scheduling in SA).
  • Google Team 1’s follow-up [5] shows (in Figure 7) that pre-training improves results only when pre-training on essentially the same design. Google could be using RL when performing multiple revisions to IC designs — a valid context, but not covered in [1]. Besides, commercial EDA tools are orders of magnitude faster (running from scratch) than RL (Table 3), so pre-training RL does not close the gap with [1].
@@
    RL’s placement of macros on a coarse grid limits their locations (Figure 1). When a human designer ignored the coarse grid, they found better macro locations [8]. Commercial EDA tools also avoid this limitation and outperform Google CT/RL.
@@ 8 Conclusions
Electronic Design Automation has been relying on common AI methods for dozens of years, including A*-search for wire routing, SAT-solving and inductive theorem-provers for verification, etc. Attempts to use ML for combinatorial optimization are more recent. This meta-analysis discusses the reproduction and evaluation of results in the Nature paper [1] on ML for macro placement, as well as the validity of methods, results and claims in that paper. In [1], we find a smorgasboard of questionable practices in ML [87] including irreproducible research practices, multiple variants of cherrypicking, misreporting, and likely data contamination (leakage). Based on crosschecked newer data, we draw conclusions with ample redundancy (resistant to isolated mistakes): the integrity of [1] is substantially undermined owing to errors in the conduct, analysis and reporting of its study. Omissions, inconsistencies, mistakes, and misrepresentations impacted methods, data, results and interpretation in [1]. Table 4 makes it clear that serious issues were raised at Google about the work many months before the Nature publication [1].
@@ 8.1 Conclusions about [1]
We crosscheck the results reported in [6, 7, 8] and also account for [2, 4, 5, 52], then summarize conclusions drawn from these works. This confirms many of the initial doubts about [1] and identifies additional deficiencies. As a result, it is clear that [1] is misleading in several ways, such that the readers can have no confidence in the top-line claims of [1]. nor its conclusions. [1] did not improve SOTA while the methods and results of the original paper were not reproducible from the descriptions provided, contrary to stated editorial policies at Nature (see Section 8.3). The reliance on proprietary TPU designs for evaluation, along with insufficient reporting of experiments, continues to obstruct reproducibility of the methods and the results. Google Team 2 [6] had access to Google internal code whereas the UCSD Team [8] reverse-engineered and/or reimplemented missing components. Google Team 2 and the UCSD team drew consistent conclusions from similar experiments, and each team made additional observations.
@@ 8.2 Conclusions for chip-design tech
    Machine-learning from entire chip designs is hard: learning from diverse designs might only improve runtime and not quality in typical cases [5]. Learning from earlier versions of the same design can be useful in some cases, but should be compared to warm-starting Simulated Annealing with an initial placement and adaptive temperature schedule.
@@
  6. 6.
    A 2024 effort from China [89] compared seven techniques for mixed-size placement using their new independent evaluation framework with 20 circuits (7 with macros). End-to-end results for chip metrics show that post-[1] ML-based techniques lag behind RePlAce [45] (embedded in OpenROAD) and other optimization-based techniques: DREAMPlace (a GPU-based variant of the RePlAce algorithm) [51] and AutoDMP (a Bayesian Optimization wrapper around DREAMPlace) [65]. Despite the obvous need to replicate the methods from [1], the authors of [89] were unable to provide such results.
@@ 8.3 Policy implications
Theoretical arguments and empirical evidence suggest that numerous published papers across various fields cannot be replicated and are likely false [10, 9, 41, 71, 77, 85]. The developments with [1] add to the so-called reproducibility crisis that undermines trust in published research results [72, 77]. In response to this crisis, some observers say “reproduce or it didn’t happen” [81].
Retraction Watch now tracks 5000 retractions per year, including prominent cases of research misconduct [75, 77]. Per [72], “research misconduct is a serious problem and (probably) getting worse”, which makes it even more important to separate honest mistakes from deliberate exaggerations and misconduct [73, 74, 75, 78, 85]. To this end, see Table 4. Institutional response is needed, and opportunities for future reforms are discussed in [71, 78, 82, 86]. Particularly important are faster, more numerous retractions [60] and clarity in Nature retraction notices [88]. Here we make a modest contribution to this far-reaching discussion by making specific suggestions.
@@
It is unclear why Google did not allow publishing [6] (coauthored by the author of this meta-analysis), even after its results and conclusions were corroborated by the published paper [8] written at UCSD with lengthy involvement from Google. Hesitation to act on [1] is understandable, even after [6] and [7, 8] found major flaws in [1], but “a commitment to open inquiry, intellectual rigor, integrity, and collaboration” must protect legitimate research in [6].
@@
Specifically for [1], the Nature editorial [3] insisted that “the technical expertise must be shared widely.” But when manuscript authors neglect requests for public benchmarking and obstruct reproducibility, their technical claims should be viewed with suspicion [66] (especially if they later disagree with comparisons to their work [62]). Per the peer review file [2], the acceptance of the Nature paper was conditional on the release of code and data in the second revision, but this did not happen when the paper was published or later, per [8]. The Nature paper [1] was amended by the authors to claim that the code had been made available (see the “Data and Code Availability” disclaimer). But serious omissions remain in the released code. This is particularly concerning because ($i$) [1] omitted key comparisons and details, and ($ii$) fraud was alleged under oath in a California court by a Google whistleblower tasked with evaluating the project [61]. This makes reproducibility more critical.
@@ 8.4 Nature editors investigate
A year later (late September 2024), the Editor’s note was removed from the Nature article, but an authors’ addendum appeared. The addendum largely repeats the arguments from an earlier statement [62] which we discussed in Section 6, so there was little for us to modify in the present article: none of the major concerns about the Nature paper have been addressed. In particular, ”results” on one additional proprietary TPU block with undisclosed statistics do not support any serious conclusions. This only aggravates concerns about cherry picking and misreporting. The release of a pretrained model without information about training data aggravates concerns about data contamination - any public benchmark could have been used in training. We do not comment on the Google blog post,161616 https://deepmind.google/discover/blog/how-alphachip-transformed-computer-chip-design/ except that it repeats the demonstrably false claim of a full source code release that allows one to reproduce the results in the Nature paper. Among other pieces, source code for simulated annealing is still missing, and additionally the Nature results cannot be reproduced without proprietary training data and test data.